OpenStudy (anonymous):

Simplify the expression -6 × 12 + 14 − 3. Select one of the options below as your answer: A. -61 B. -58 C. -51 D. -30

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

A is not a number :)

OpenStudy (anonymous):

You can type these into a calculator to find the answer XD. Try "Wolframalpha.com", you can get the answer more easily without waiting for a reply. Touche Amistre.

OpenStudy (anonymous):

ohkay thanks

OpenStudy (anonymous):

-6 * 12 + 14-3 = -72+14-3=-72+11=-61 Hence, (A)=-61 is the answer.